OpenStudy (anonymous):

The volume of two similar solids is 1331 m3 and 729 m3. The surface area of the larger solid is 605 m2. What is the surface area, in square meters, of the smaller solid?

OpenStudy (anonymous):

121 305 405 81

OpenStudy (anonymous):

The King George

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

proportions?

OpenStudy (kinggeorge):

Definitely proportions, but in a certain way. Notice that the volume of two solids is proportional to the cube of the side lengths, while the surface area is proportional to the square of the side lengths.

OpenStudy (anonymous):

hmm

OpenStudy (kinggeorge):

So what I would do is find the proportion for the side lengths, and then square it to find the proportion of the surface area. Once you have that, you can simply plug 605m\(^2\) into the proportion.

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

well

OpenStudy (anonymous):

the cube root of 1331 is 11 while the cube root of 729 is 9

OpenStudy (anonymous):

does that help any?

OpenStudy (kinggeorge):

That's great so far. The ratio of the side lengths would then be 11/9. So the ratio of the surface areas is...?

OpenStudy (anonymous):

605/x

OpenStudy (anonymous):

I got 495

OpenStudy (anonymous):

thats not an answer choice , though :(

OpenStudy (kinggeorge):

I don't think you used the correct proportions there. If the proportion for side length is 11/9, and the proportion for surface area is the square of the proportion for side length, what is that proportion?

OpenStudy (anonymous):

11/605?

OpenStudy (anonymous):

605/11=55

OpenStudy (kinggeorge):

Not quite, sorry if I wasn't very clear with I said. It should be\[\left(\frac{11}{9}\right)^2=\frac{121}{81}\]

OpenStudy (anonymous):

AHH

OpenStudy (anonymous):

THE AREA

OpenStudy (anonymous):

121/81=605/x

OpenStudy (anonymous):

405

OpenStudy (kinggeorge):

Precisely.

OpenStudy (anonymous):

next problem

OpenStudy (anonymous):

9^3/11^3 so the area ration will be 9^2/11^2 9^2/11^2*605=405